A student is allowed to select at most n-blocks from a collection of (2n + 1) books. If the total number of ways in which he can select a book is 63, find the value of n.

Grade:11

1 Answers

Chetan Mandayam Nayakar
11 years ago

Σ2n+1Cr,(from r=1 to r=n)=(22n+1/2)-1=22n-1=63,n=3